Web Analytics

two card decks are drawn from a standard 52-card deck, without replacement. Find the probability that they are both aces.
A. 1/2652
B. 3/676
C. 1/221
D. 1/169

Verified Answer :

I'm sorry, when I answered this I'm pretty sure I wasn't in algebra two yet. The correct answer is 4/52 * 3/51 = 1/221

Answer:

Option C. 1/221

Step-by-step explanation:

Total cards in a standard deck of cards = 52

Number of ace in a deck of cards = 4

So probability to draw an ace  \(P_{1}=\frac{4}{52}\)

Since the cards were drawn without replacement means remaining cards in deck of cards now = 51

Number of aces remaining = 3

So probability of drawing another ace  \(P_{2}=\frac{3}{51}\)

Now probability to draw both the aces

= \(P_{1}\times P_{2}\)

\(\frac{4}{52}\times \frac{3}{51}=\frac{1}{13}\times \frac{1}{17}\)

= \(\frac{1}{221}\)

Option C. is the answer.